# Rational Motions of Minimal Quaternionic Degree with Prescribed Plane Trajectories

### v1

@@ Abstract
kinematics , dual quaternion , rational kinematic torse , quaternionic polynomial
@@ 1 Introduction
In this paper we consider rational rigid body motions of Euclidean three-space with a prescribed rational plane trajectory (a rational torse). We provide answers to the following questions:
@@
While point trajectories in kinematics have been investigated extensively, much less in known for the trajectory of planes (and of lines). A standard reference on theoretical kinematics features a short section on that topic [7, Section VI.7]. The paper [8] studies rigid body transformations of points, lines, and planes in a Clifford algebra setting and discusses examples from computer vision and geometric constraint solving. An additional motivation to study the kinematics of plane trajectories is the close relation between torses and developable surface [9].
@@ 2.3 Rational Curves, Torses, and Motions
We use square brackets around $\gamma$, $u$, and $C$ to denote points of the respective projective spaces that these entities lie in. It is clear from Definition 1 that rational curves, torses and motions can be represented not just by rational functions but also by polynomials in the indeterminate $t$ and with coefficients from $\mathbb{D}\mathbb{H}$. This is what we will usually do. If $[C]$ is a polynomial representation of a rational motion, we call $C$ a motion polynomial, cf. [11]. If $[u]$ is a polynomial representation of a rational torse, we call $u$ a plane polynomial. The extensions of the actions (3) and (5) from dual quaternions to motion polynomials provides polynomial descriptions for rational curves and torses, respectively. Note that we also constant points, plans, and rigid body displacements are to be considered as points of the respective projective spaces.
@@ 2.4 Problem Statement and Simplifications
Given a rational torse $[u]=[\vec{u}+\varepsilon u_{0}]=[u_{1}\mathbf{i}+u_{2}\mathbf{j}+u_{3}\mathbf{k}+\varepsilon u_{0}]$, we are looking for a rational motion $[C]=[P+\varepsilon D]$ and a constant plane $[w]$ (the “moving plane”) such that $[u]=[{C}_{\varepsilon}w{C}^{\ast}]$ holds. For any plane $[w]$ there exists a suitable rigid body displacement $a$ such that ${a}_{\varepsilon}v{a}^{\ast}=\mathbf{k}$. The dual quaternion $a$ may be absorbed in $[C]$ (replace $C$ by $Ca$) so that we can assume without loss of generality that the moving plane is $[\mathbf{k}]$. We thus will study the equation
@@
The rational torse $[u]$ is given, the rational motion $[C]$ is sought. As we shall see (Theorem 1), solutions to (6) exist if and only if a natural assumption on $[u]$, rationality of its Gauss map, is satisfied. If a solution exists, it is not unique: We can always multiply $[C]$ from the right with a rational motion $[E]$ that fixes the plane $[\mathbf{k}]$. The motion $[E]$ is then a planar motion, that is, a motion in $\operatorname{SE}(2)$, that necessarily is of the shape $E=e_{0}+e_{3}\mathbf{k}+\varepsilon(e_{5}\mathbf{i}+e_{6}\mathbf{j})$. Indeed, ${E}_{\varepsilon}\mathbf{k}{E}^{\ast}=(e_{0}^{2}+e_{3}^{2})\mathbf{k}$ whence $[{E}_{\varepsilon}\mathbf{k}{E}^{\ast}]=[\mathbf{k}]$. Given the non-uniqueness of solutions to Equation (6), it is natural to search for “simple” solutions and it is equally natural to define simplicity in terms of the motion degree:
@@
Call a polynomial in $w=w_{0}+w_{1}\mathbf{i}+w_{2}\mathbf{j}+w_{3}\mathbf{k}+\varepsilon(w_{4}+w_{5}\mathbf{i}+w_{6}\mathbf{j}+w_{7}\mathbf{k})\in\mathbb{D}\mathbb{H}[t]$ reduced if it has no real polynomial factor of positive degree. This can be expressed using the $\gcd$ of real polynomials. For $w\in\mathbb{D}\mathbb{H}[t]$, define its real greatest common divisor as $\operatorname{realgcd}(w)\coloneqq\gcd(w_{0},w_{1},\ldots,w_{7})$. Now, $w$ is reduced if and only if $\operatorname{realgcd}(w)=1$. The degree of a rational curve, torse, or motion $[x]$ equals $\deg x$ if and only if $x$ is reduced.
@@
with some $\alpha$, $\beta$, $\gamma$, $\delta\in\mathbb{R}$ such that $\alpha\delta-\beta\gamma\neq 0$ and multiply away denominators. An alternative would be the use of homogeneous polynomials throughout this text. We don’t do this because of the overhead in notation and also existing literature on motion polynomials avoids this.
@@ 3.1 The Primal Part
In what follows, we embed the complex number field $\mathbb{C}$ into $\mathbb{H}$ as sub-algebra generated by $1$ and $\mathbf{k}$.
@@ 3.2 The Dual Part
$d_{0}=\frac{v_{2}d_{1}-v_{1}d_{2}-q_{3}u_{0}}{2(q_{0}^{2}+q_{3}^{2})},\quad d_{3}=-\frac{v_{1}d_{1}+v_{2}d_{2}-q_{0}u_{0}}{2(q_{0}^{2}+q_{3}^{2})}$ (14)
@@
$d_{1}=-\tfrac{1}{2}u_{0}v_{3}^{-1}q_{2},\quad d_{2}=\tfrac{1}{2}u_{0}v_{3}^{-1}q_{1}$ (16)
@@
$d_{1}=-\tfrac{1}{2}u_{0}v_{3}^{-1}q_{2}+\lambda q_{3},\quad d_{2}=+\tfrac{1}{2}u_{0}v_{3}^{-1}q_{1}-\lambda q_{0}$
for arbitrary $\lambda\in R$. By Lemma 4 below, with $p=q_{0}^{2}+q_{3}^{3}$, $x=d_{1}$, $y=d_{2}$, $a=q_{3}$ and $b=-q_{0}$, there exists $\lambda\in R$ such that the desired degree bound can be achieved. We still need to argue that the conditions to apply Lemma 4 are met. These are
@@ 4 Motions of Minimal Degree
Remark 3.
Only kinematic torses can arise as trajectories of a plane under a rational motion. They can be characterized among all rational torses as those having a rational Gauss map $n\coloneqq\vec{u}/\sqrt{N(u)}$. Thus, the tangent planes of a cone of revolution form the planes of a kinematic rational torse while the tangent planes of a general quadratic cone don’t. The former obviously arise trajectory of a rational motion, the rotation around the cone’s axis. For the latter, Equation (7) and also Equation (8) have no solution.
The following theorem contains the main results of this article. It uses the notion of “essential uniqueness” of the rational motion $[C]$. By this we mean that $[C]$ is unique up to right multiplication with a constant displacement $[e]$ that fixes the plane $[\mathbf{k}]$, that is $[{e}_{\varepsilon}\mathbf{k}{e}^{\ast}]=[\mathbf{k}]$.
@@
    There exists a rational motion $[C]$ of degree $\deg u-\frac{1}{2}\deg n$ with trajectory $[u]$.
@@
    All rational motions with trajectory $[u]$ are obtained by composing $[C]$ from the right with any rational motion $[E]$ that fixes the plane $[\mathbf{k}]$. [^2] This implies that $E$ is given by a polynomial of the shape $E=e_{0}+e_{3}\mathbf{k}+\varepsilon(e_{5}\mathbf{i}+e_{6}\mathbf{j})$ with $e_{0}$, $e_{3}$, $e_{5}$, $e_{6}\in\mathbb{R}[t]$.
@@
Finally, any solution $[\hat{C}]$ can be reduced to the minimal degree solution $[C]$ by splitting of suitable motion polynomial right factors $E$ as shown in Lemmas 6 and 7. Thus, $[\hat{C}]$ is conversely obtained from $[C]$ by composition with $[E]$ from the right. ∎
In order to actually compute rational motions of minimal degree we may proceed as in Examples 4 and 5. Figure 1 visualizes the trajectories of the plane $[\mathbf{k}]$ for the rational motion $[C]$ given by (18) in Example 4 and the rational motion $[\hat{C}]$ given by (20) in Example 5. The positions of the moving plane are represented by moving rectangles in both cases. Two corresponding rectangles lie in the same plane. In Figure 1, a pair of corresponding points is connected by dotted lines. The orientation of the rectangular frames relative to this connecting line changes and also their lengths vary. Thus, the two motions are indeed essentially different.
Figure 1: Two different motions with identical plane trajectory.
@@ 5 Conclusion and Future Research
A natural next question is to study the problem of creating a given rational ruled surface as trajectory of a straight line undergoing a rational motion. Again, there is the necessary condition on rationality of the ruled surface’s spherical image. Apart from that, little seems to be known. We expect, however, that techniques similar to the ones used by us in this article will provide a complete solution.
Once torses or ruled surfaces are described as plane or line trajectories of rational motions of low degree, motion factorization techniques can be used for the creation of mechanical linkages to “draw” them. It seems natural to adapt the techniques of [5, 6] to the case of plane and line trajectories, thus extending Kempe’s Universality Theorem to kinematic rational torses and to ruled surfaces. This is not an automatic task as one important insight of this article is that the rational motions to create a given line trajectory or a given plane trajectory have quite different characteristics.
@@
A further and rather different line of research starts with the fundamental Equation (7) that we solved for an undetermined motion polynomial $C$ and a real polynomial $h$. It can be viewed as a generalization of Equation (8), an equation that describes the quaternionic pre-image of a polynomial Pythagorean-hodograph curve [17]. It can also be regarded as a special instance of an equation of Sylvester type. The Sylvester equation has been studied extensively in diverse algebraic context and in particular in the context of quaternionic function theory. However, in contrast to [22], we are not interested in “slice semi-regular quaternionic solution functions” to the general Sylvester equation $ax+xb=c$ but in low-degree polynomial solutions. We suggest the study of low degree polynomial solutions to Sylvester type equations in a general algebraic context might an interesting topic of future research.

### v3

@@ Abstract
kinematics , dual quaternion , rational kinematic torse , developable surface , quaternionic polynomial
@@ 1 Introduction
In this paper we consider rational rigid body motions of Euclidean three-space with a prescribed rational plane trajectory (a rational torse or rational developable surface). [^1] We prefer the term “torse” over “developable surface” as it emphasizes that we consider one-parametric sets of planes (curves in dual space), not their envelopes (developable surfaces). We provide answers to the following questions:
@@
While point trajectories in kinematics have been investigated extensively, much less in known for the trajectory of planes (and of lines). A standard reference on theoretical kinematics features a short section on that topic [7, Section VI.7]. The paper [8] studies rigid body transformations of points, lines, and planes in a Clifford algebra setting and discusses examples from computer vision and geometric constraint solving. An additional motivation to study the kinematics of plane trajectories is the close relation between torses and developable surfaces [9].
@@ 2.3 Rational Curves, Torses, and Motions
[…]
@@ 2.4 Problem Statement and Simplifications
[…]
@@
[…]
@@
[…]
@@
with some $\alpha$, $\beta$, $\gamma$, $\delta\in\mathbb{R}$ such that $\alpha\delta-\beta\gamma\neq 0$ and multiply away denominators. An alternative would be the use of homogeneous polynomials throughout this text. We do not do this because of the overhead in notation and also existing literature on motion polynomials avoids this.
@@ 3.1 The Primal Part
In what follows, we embed the complex number field $\mathbb{C}$ into $\mathbb{H}$ as sub-algebra generated by 111 and $\mathbf{k}$.
@@ 3.2 The Dual Part
$d_{0}=\frac{v_{2}d_{1}-v_{1}d_{2}-q_{3}w_{0}}{2(q_{0}^{2}+q_{3}^{2})},\quad d_{3}=-\frac{v_{1}d_{1}+v_{2}d_{2}-q_{0}w_{0}}{2(q_{0}^{2}+q_{3}^{2})}$ (14)
@@
$d_{1}=-\tfrac{1}{2}w_{0}v_{3}^{-1}q_{2},\quad d_{2}=\tfrac{1}{2}w_{0}v_{3}^{-1}q_{1}$ (16)
@@
$d_{1}=-\tfrac{1}{2}w_{0}v_{3}^{-1}q_{2}+\lambda q_{3},\quad d_{2}=+\tfrac{1}{2}w_{0}v_{3}^{-1}q_{1}-\lambda q_{0}$
for arbitrary $\lambda\in R$. By Lemma 4 below, with $p=q_{0}^{2}+q_{3}^{2}$, $x=d_{1}$, $y=d_{2}$, $a=q_{3}$ and $b=-q_{0}$, there exists $\lambda\in R$ such that the desired degree bound can be achieved. We still need to argue that the conditions to apply Lemma 4 are met. These are
@@ 4 Motions of Minimal Degree
Remark.
Only kinematic torses can arise as trajectories of a plane under a rational motion. They can be characterized among all rational torses as those having a rational Gauss map $n\coloneqq\vec{u}/\sqrt{N(u)}$. Thus, the tangent planes of a cone of revolution form the planes of a kinematic rational torse while the tangent planes of a general quadratic cone do not. The former obviously arise trajectory of a rational motion, the rotation around the cone’s axis. For the latter, Equation (7) and also Equation (8) have no solution.
[…]
@@
[…]
@@
[…]
@@
[…]
Figure 1: Two different minimal degree motions with identical plane trajectory.
Example 7.
As a concluding example, lets consider the Darboux motion, the famous rational space motion with only planar trajectories [7, Ch. 9, § 3]. It can be thought of as the composition of a Cardan motion in a horizontal plane with a harmonic oscillation along a vertical axis. Figure 2 visualizes the motion of three planes undergoing a Darboux motion. The first trajectory is that of a generic plane where both, the torse and the Gauss map are of degree two. Hence, the Darboux motion is minimal for this torse. The second trajectory is that of a horizontal plane. We clearly observe that the planes’ normals have a fixed direction whence the degree of the Gauss map is zero. The Darboux motion is minimal but not unique. It can be composed from the right with a suitable translational motion with zero vertical component. This is illustrated by the smaller frames. Clearly, they lie in the same plane but they also change their relative position with respect to the larger frame.
Figure 2: Plane trajectories of the Darboux motion.
@@ 5 Conclusion and Future Research
A natural next question is to study the problem of creating a given rational ruled surface as trajectory of a straight line undergoing a rational motion. As demonstrated in the pre-print [20], fundamental statements are quite similar to the torse case but proof details differ.